OpenStudy (anonymous):

You invest $500 in an account with an annual interest rate of 1.1%, compounded continuously. How much money is in the account after 5 years? Round your answer to the nearest whole number. A: $528 B: $530 C: $545 D: $560

OpenStudy (anonymous):

Using the equation: A=Pe^(rt) A is the Amount after the time or your answer. P is the principal amount in this case the 500 r is the interest rate so 1.1% plug it in as .011 t is the time in years so 5 years

OpenStudy (anonymous):

Can't you just give me the answer? Because this isn't making any sense to me.

jimthompson5910 (jim_thompson5910):

A=Pe^(rt) A=500*2.718^(0.011*5) A = ???

jimthompson5910 (jim_thompson5910):

let me know what you get

OpenStudy (anonymous):

2,527 is what I got.

OpenStudy (anonymous):

Which is wrong...

OpenStudy (anonymous):

Oh, I did it again and got 258. Thanks for your help!

OpenStudy (anonymous):

Sorry - 528

jimthompson5910 (jim_thompson5910):

you are correct, nice work